# Symmetries of the periodic Fredkin chain

### v1

@@ 3.3. Proof of Theorem 2
[…]
Formula (3.9) is simply (2.14) written in terms of the Pauli spin operators; (3.10) and (3.11) are analogues or (3.2).
Now, one can prove the result of action of the operators $\Sigma^{\pm}$ at the vectors $\left\lvert v_{S^{z}}\right\rangle$. Using the residue formulas (3.2) and (3.10), we find
@@
Let us now consider commutativity of the operators $\Sigma^{\pm}$ with the Hamiltonian, relation (3.4). To this end it is useful to consider more general operators:
@@
Here, as above, $\sigma_{i}^{0}\equiv 1$, $\sigma_{i}^{\pm 1}\equiv\sigma_{i}^{\pm}$, and the subscripts $i_{1},\dots,i_{n}$, indicate copies of $\mathbb{C}^{2}$ where the Pauli spin operators act, $1\leqslant i_{1}<\dots<i_{n}\leqslant N$. The superscript $Q$ can be regarded as a “charge” of the operator. Note that with this notation $\Sigma^{\pm}\equiv\Sigma^{\pm 1}_{1,\dots,N}$, and we will follow the convention to drop the subscripts when $n=N$, i.e., writing $\Sigma^{Q}\equiv\Sigma^{Q}_{1,\dots,N}$. The following observation is crucial.
@@
$[\Sigma^{Q},\mathrm{F}_{i,i+1,i+2}]=0,\qquad Q=0,\pm 1,\dots,\pm N.$ (3.17)
@@
Note that this decomposition simply follows from the core definition (3.16) and it can also be derived formally by writing a residue formula by analogue with (3.2), and next evaluating the residues after the summation. The decomposition (3.18) implies that (3.17) will follow if $[\Sigma_{123}^{Q},\mathrm{F}_{123}]=0$ for $Q=0,\pm 1,\pm 2,\pm 3$. Explicitly, the operators appearing here are
@@ 4. Symmetry algebra
Here, we focus on further discussion of properties of the operators $\Sigma^{\pm}$, such as which algebra they span and what is a role of the operator $\mathrm{S}^{z}$ in this construction. We start with comments about one more operator which exists in the case of $N$ even and next proceed to the results.
@@ 4.1. Preliminaries
In the case of $N$ even, one can introduce one more operator, $\Xi$, which when acting at the ground states maps the states $\left\lvert v_{0}\right\rangle\equiv\left\lvert v_{0}^{+}\right\rangle$ and $\left\lvert v_{0}^{-}\right\rangle$ to each other,
@@
As it is clear from (4.3), the operator $\Xi$ is represented by a diagonal matrix with 00’s, $1$’s and $-1$’s in its entries, with 00’s responsible for cutting out all sectors with $S^{z}=\pm 1,\dots,\pm N$, whereas in the sector with $S^{z}=0$ it has only $1$’s and $-1$’s in an equal number. In other words, $\Xi$ acts nontrivially only in the $S^{z}=0$ sector of the whole Hilbert space of the Hamiltonian. It commutes with the Hamiltonian and anti-commutes with the cyclic shift operator:
$[\Xi,\mathrm{H}]=0,\qquad\mathrm{C}\Xi\mathrm{C}^{-1}=-\Xi.$ (4.4)
The first relation may require a proof though we find it satisfied in all checks; the second one is obvious from (4.3). These relations mean that $\Xi$ maps every state in the $S^{z}=0$ sector to another state in this sector with the same eigenvalue of $\mathrm{H}$, but alternated eigenvalue of the cyclic shift operator, $C\mapsto-C$. Existence of such an unusual operator may explain a clear visible (in numerical checks) double degeneracy of the whole spectrum of the Hamiltonian. It complements, for the case of $S^{z}=0$, doubling of states in the $S^{z}\neq 0$ sectors following from the spin-reversal symmetry.
@@ 4.2. The Lie algebra of raising and lowering operators
Now we focus on establishing an algebra the operators $\Sigma^{\pm}$ generate. We denote this algebra by $\mathfrak{g}$. Our findings can be summarized as follows.
@@
As in [P-25], to formulate the above conjecture we have performed the following calculation. In addition to $\Sigma^{\pm}$, we define $\Sigma^{z}=[\Sigma^{+},\Sigma^{-}]$ and construct triples $\Lambda^{\pm,z}$, $\Phi^{\pm,z}$, etc, introducing them recursively, first $\Lambda^{\pm}=\pm\left[\Sigma^{z},\Sigma^{\pm}\right]$, $\Lambda^{z}=\left[\Lambda^{+},\Lambda^{-}\right]$, next $\Phi^{\pm}=\pm\left[\Lambda^{z},\Lambda^{\pm}\right]$, $\Phi^{z}=\left[\Phi^{+},\Phi^{-}\right]$, and so on. An important observation which can be easily done at this stage is that the operators $\Sigma^{z},\Lambda^{z},\Phi^{z},\dots$ form an abelian algebra, and therefore they can considered as linear combinations of the Cartan subalgebra elements $h_{i}$. Hence, the rank of the algebra can be identified by how many operators $\Sigma^{z},\Lambda^{z},\Phi^{z},\dots$ are linearly independent.
@@
where the number of terms is equal to the rank. Solving relations $[e_{i},f_{j}]=0$, $i\neq j$, for the coefficients $a_{i},b_{i},\dots$ one finds a unique solution up to a permutation of triples of Chevalley generators. The solution is fixed by finding $\rho_{i}$ from the relations $[h_{i},e_{i}]=2e_{i}$, that corresponds to the canonical normalization of the Cartan matrix with $A_{ii}=2$.
At the last step, the off-diagonal entries of the Cartan matrix are computed according to the relations $[h_{i},e_{j}]=A_{ij}e_{j}$. A particular order among the triples of Chevalley generators should be fixed, such that the Cartan matrix can be identified according to the well-known classification by Dynkin diagrams. In our problem we meet Cartan matrices for the Lie algebras of $B$- and $C$-types. Recall that the off-diagonal nonzero entries of the Cartan matrix, in the case of $B$-type algebra are
@@ 4.3. A central element
Noting that $[\mathrm{S}^{z},\Sigma^{\pm}]=\pm\Sigma^{\pm}$ and one can expect the operator $\mathrm{S}^{z}$ can be expressed terms of the Cartan subalgebra elements. It turns out that such a construction involves one more operator which appears to be a central element with respect to the generators of $\mathfrak{g}$. This can be formulated as follows.
@@
There exists a nontrivial operator $p$ which is a central element with the respect to the algebra $\mathfrak{g}$, and it has the form
@@
We call $p$ central element since it commutes with the operators $\Sigma^{\pm}$ and hence with all the Chevalley generators, $[p,e_{i}]=[p,f_{i}]=[p,h_{i}]$.
Formula (4.10) can be viewed as expression for the operator $\mathrm{S}^{z}$ in terms of elements of $\mathfrak{g}$ and the operator $p$. Examples up to $N=8$ read
$\begin{gathered}\alpha_{1}=2,\quad\alpha_{2}=\frac{3}{2},\qquad N=3,4,\\ \alpha_{1}=\frac{5}{2},\quad\alpha_{2}=4,\quad\alpha_{3}=\frac{9}{2},\qquad N=5,\\ \alpha_{1}=3,\quad\alpha_{2}=5,\quad\alpha_{3}=3,\qquad N=6,\\ \alpha_{1}=\frac{7}{2},\quad\alpha_{2}=6,\quad\alpha_{3}=\frac{15}{2},\quad\alpha_{4}=8,\qquad N=7,\\ \alpha_{1}=4,\quad\alpha_{2}=7,\quad\alpha_{3}=9,\quad\alpha_{4}=5,\qquad N=8.\end{gathered}$ (4.11)
Coincidence of the coefficients in the cases of $N=3$ and $N=4$ can be explained by appearance of the same algebra ($B_{2}=C_{2}$) in these two cases. Furthermore, an interesting conclusion which can be made by inspecting these examples, is that in the case of $C$-type Lie algebra ($N$ even) formula (4.10) exactly coincides with that obtained in [P-25] in the context of the periodic Motzkin chain. This suggests, that the relation (4.10) is probably independent of the choice of the initial representation of the $\mathfrak{sl}_{2}$ algebra of local spin operators.
Inspecting (4.11) one can make some further conjectures. For example, one can guess that $\alpha_{1}=N/2$ for $N>3$, and $\alpha_{2}=N-1$ for $N>4$. However, to get more insight on these numbers, a more output is needed for further values of $N$. This is not a quite simple task since it requires an explicit construction of the Chevalley generators in each particular case, to establish the corresponding relation (4.10).
@@ 5. Conclusion
In this paper, we have considered the Fredkin spin chain with periodic boundary conditions basing on the result of Salberger and Korepin [SK-17, SK-18] about explicit form of the ground states (Theorem 1).
We have showed that there exist raising and lowering operators which maps these states to each other and which commute with the Hamiltonian (Theorem 2). We have also provided several conjectures about other properties of these operators. Conjecture 1 says that these operators actually act non-trivially only in the subspace of cyclic invariant states. Conjecture 2 shows that these operators can be (somewhat unexpectedly) be expressed in terms of the non-diagonal components of total spin operator. Conjecture 3 states that they span in $B$- or $C$-type Lie algebra, according to the parity of number of sites of the chain. Conjecture 4 asserts that the symmetry algebra of the model is actually wider and extended by a central element contained in the third component of total spin operator.
In our previous study devoted to the Motzkin chain [P-25] the result analogous to Theorem 2 was formulated as a conjecture. In fact, the proof we give here for Theorem 2 can be directly generalized to the spin-$1$ case of Motzkin chain. This is intimately related to fact that the lowering an raising operators have essentially the same structure for both models.
Conjecture 1 can also be formulated in the context of Motzkin chain. It is not however clear how it can proven in either case. Concerning Conjecture 2, it seems plausible that it can be proven be suitably applying the technique exposed in [ZF-11]. Conjectures 3 and 4 are direct analogues of the corresponding conjectures made in the context of Motzkin chain in [P-25]. We have observed that the central element for the Fredkin chain with even number of sites appears to be related to the third component of total spin operator by exactly the same relation as it was found in the context of the Motzkin chain. In both cases the symmetry algebra is the $C$-type Lie algebra of the same rank. It would be interesting to understand group theoretical backgrounds of these results. In addition, the coefficients $\gamma_{i}$ and $\alpha_{i}$ appearing in Conjecture 2 and Conjecture 4, respectively, certainly have some combinatorial interpretations that can also be a subject of further study.
@@ Acknowledgments
The author is indebted for stimulating discussions to Nikolai M. Bogoliubov, Vladimir E. Korepin, and Vitaly O. Tarasov.

### v2

@@ 3.3. Proof of Theorem 2
Here, $\left\lvert\Uparrow\right\rangle$ is the state “all spins up”, $\left\lvert\Uparrow\right\rangle=\left\lvert\uparrow\right\rangle^{\otimes N}$.
Formula (3.8) is simply (2.14) written in terms of the Pauli spin operators; (3.9) and (3.10) are analogues of (3.2).
Now, one can prove the result of action of the operators $\Sigma^{\pm}$ on the vectors $\left\lvert v_{S^{z}}\right\rangle$. Using the residue formulas (3.2) and (3.9), we find
@@
Let us now consider commutativity of the operators $\Sigma^{\pm}$ with the Hamiltonian, relation (3.4). It is useful to consider more general operators:
@@
Here, as above, $\sigma_{i}^{0}\equiv 1$, $\sigma_{i}^{\pm 1}\equiv\sigma_{i}^{\pm}$, and the subscripts $i_{1},\dots,i_{n}$, indicate copies of $\mathbb{C}^{2}$ where the Pauli spin operators act, $1\leqslant i_{1}<\dots<i_{n}\leqslant N$. The superscript $Q$ can be interpreted as a “charge” of the operator. Note that with this notation $\Sigma^{\pm}\equiv\Sigma^{\pm 1}_{1,\dots,N}$, and below we drop the subscripts when $n=N$, i.e., writing $\Sigma^{Q}\equiv\Sigma^{Q}_{1,\dots,N}$. The following observation is crucial.
@@
$\Sigma^{Q},\mathrm{F}_{i,i+1,i+2}]=0,\qquad Q=0,\pm 1,\dots,\pm N.$ (3.16)
@@
Note that this decomposition simply follows from the core definition (3.15) and it can also be derived formally by writing a residue formula by analogue with (3.2), and next evaluating the residues after the summation. The decomposition (3.17) implies that (3.16) will follow if $\Sigma_{123}^{Q},\mathrm{F}_{123}]=0$ for $Q=0,\pm 1,\pm 2,\pm 3$. Explicitly, the operators appearing here are
@@ 4. Symmetry algebra
Here, we focus on further discussion of properties of the operators $\Sigma^{\pm}$, such as which algebra they span and what the role of the operator $\mathrm{S}^{z}$ is in this construction. We start with comments about one more operator which exists in the case of $N$ even and next proceed to the results.
@@ 4.1. Preliminaries
In the case of even $N$, one can introduce one more operator, $\Xi$, which, when acting on the ground states maps the states $\left\lvert v_{0}\right\rangle\equiv\left\lvert v_{0}^{+}\right\rangle$ and $\left\lvert v_{0}^{-}\right\rangle$ to each other,
@@
As it is clear from (4.3), the operator $\Xi$ is represented by a diagonal matrix with 0’s, 11’s and 1-1’s in its entries, with 0’s responsible for cutting out all sectors with $S^{z}=\pm 1,\dots,\pm N/2$. In the sector with $S^{z}=0$ the entries are all nonzero, represented by $1$’s and 1-1’s in an equal number. In other words, $\Xi$ acts nontrivially only in the $S^{z}=0$ sector of the whole Hilbert space of the model. It commutes with the Hamiltonian and anti-commutes with the cyclic shift operator:
$\Xi,\mathrm{H}]=0,\qquad\mathrm{C}\Xi\mathrm{C}^{-1}=-\Xi.$ (4.4)
The first relation may require a proof though we find it satisfied in all checks; the second one is obvious from (4.3). These relations mean that $\Xi$ maps every state in the $S^{z}=0$ sector to another state in this sector with the same eigenvalue of $\mathrm{H}$, but alternated eigenvalue of the cyclic shift operator, $C\mapsto-C$. Existence of such an unusual operator may explain a clearly visible (in numerical checks) double degeneracy of the whole spectrum of the Hamiltonian. It complements, for the case of $S^{z}=0$, the doubling of states in the $S^{z}\neq 0$ sectors following from the spin-reversal symmetry.
@@ 4.2. The Lie algebra of raising and lowering operators
Now we focus on establishing the algebra the operators $\Sigma^{\pm}$ generate. We denote this algebra by $\mathfrak{g}$. Our findings can be summarized as follows.
@@
As in [P-25], to formulate the above conjecture we have performed the following calculation. In addition to $\Sigma^{\pm}$, we define $\Sigma^{z}=[\Sigma^{+},\Sigma^{-}]$ and construct triples $\Lambda^{\pm,z}$, $\Phi^{\pm,z}$, etc, introducing them recursively, first $\Lambda^{\pm}=\pm\left[\Sigma^{z},\Sigma^{\pm}\right]$, $\Lambda^{z}=\left[\Lambda^{+},\Lambda^{-}\right]$, next $\Phi^{\pm}=\pm\left[\Lambda^{z},\Lambda^{\pm}\right]$, $\Phi^{z}=\left[\Phi^{+},\Phi^{-}\right]$, and so on. An important observation which can be easily done at this stage is that the operators $\Sigma^{z},\Lambda^{z},\Phi^{z},\dots$ form an abelian algebra, and therefore they can be considered as linear combinations of the Cartan subalgebra elements $h_{i}$. Hence, the rank of the algebra can be identified by how many operators $\Sigma^{z},\Lambda^{z},\Phi^{z},\dots$ are linearly independent.
@@
where the number of terms is equal to the rank. Solving relations $e_{i},f_{j}]=0$, $i\neq j$, for the coefficients $a_{i},b_{i},\dots$ one finds a unique solution up to a permutation of triples of Chevalley generators. The solution is fixed by finding $\rho_{i}$ from the relations $h_{i},e_{i}]=2e_{i}$, which correspond to the canonical normalization of the Cartan matrix with $A_{ii}=2$.
At the last step, the off-diagonal entries of the Cartan matrix are computed according to the relations $h_{i},e_{j}]=A_{ij}e_{j}$. A particular order among the triples of Chevalley generators should be fixed, such that the Cartan matrix can be identified according to the well-known classification by Dynkin diagrams. In our problem, we meet Cartan matrices for the Lie algebras of $B$- and $C$-types. Recall that the off-diagonal nonzero entries of the Cartan matrix, in the case of $B$-type algebra are
@@ 4.3. A central element
Noting that $\mathrm{S}^{z},\Sigma^{\pm}]=\pm\Sigma^{\pm}$, one can expect that the operator $\mathrm{S}^{z}$ can be expressed in terms of the Cartan subalgebra elements. It turns out that such a construction involves one more operator which appears to be a central element with respect to the generators of $\mathfrak{g}$. This can be formulated as follows.
@@
There exists a nontrivial operator $p$ which is a central element with respect to the algebra $\mathfrak{g}$, and it has the form
@@
We call $p$ a central element since it commutes with the operators $\Sigma^{\pm}$ and hence with all the Chevalley generators, $p,e_{i}]=[p,f_{i}]=[p,h_{i}]$.
Formula (4.10) can be viewed as an expression for the operator $\mathrm{S}^{z}$ in terms of elements of $\mathfrak{g}$ and the operator $p$. The coefficients $\alpha_{k}$ for $N\leqslant 10$ are given in Table 2.
Table 2. Coefficients $\alpha_{k}$
NN $\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\alpha_{4}$ $\alpha_{5}$
3 22 $\frac{3}{2}$
4 22 $\frac{3}{2}$
5 $\frac{5}{2}$ 44 $\frac{9}{2}$
6 33 55 33
7 $\frac{7}{2}$ 66 $\frac{15}{2}$ 88
8 44 77 99 55
9 $\frac{9}{2}$ 88 $\frac{21}{2}$ 1212 $\frac{25}{2}$
10 55 99 1212 14 $\frac{15}{2}$
Coincidence of the coefficients for $N=3$ and $N=4$ can be explained by the appearance of the same algebra ($B_{2}=C_{2}$) in these two cases. Furthermore, an interesting conclusion which can be made by inspecting these examples, is that in the case of $C$-type Lie algebra ($N$ even) formula (4.10) exactly coincides with that obtained in the context of the periodic Motzkin chain (of length $N/2$) [P-25]. This suggests that relation (4.10) is probably independent of the choice of the initial representation of the $\mathfrak{sl}_{2}$ algebra of local spin operators.
Inspecting values of $\alpha_{k}$’s in Table 2 one can make some further conjectures. For example, one can guess that $\alpha_{1}=N/2$ for $N\geqslant 4$, and $\alpha_{2}=N-1$ for $N\geqslant 5$. It looks also plausible that $\alpha_{3}=\frac{3}{2}N-3$ for $N\geqslant 7$.
However, to gain more insight into these numbers, outputs are needed for further values of $N$ (beyond the $N=10$ case). This is not quite a simple computational task, as it requires explicit construction of the Chevalley generators to establish the corresponding relation (4.10). Along the way, one has to solve the equations $e_{i},f_{j}]=0$, $i\neq j$, for the unknowns $a_{i},b_{i},\dots$ in (4.7). These equations are quadratic with integer coefficients that grow enormously with $N$ (e.g., in the $N=10$ case, they consist of almost two hundred digits).
@@ 5. Conclusion
In this paper, we have considered the Fredkin spin chain with periodic boundary conditions. Our starting point is the result of Salberger and Korepin [SK-17, SK-18] about the explicit form of the ground states (Theorem 1). The results are summarized in one theorem and four conjectures. The latter have been formulated by studying the chain of length $N\leqslant 10$.
We have shown that there exist raising and lowering operators which map these states to each other and which commute with the Hamiltonian (Theorem 2). We have also provided several conjectures about other properties of these operators. Conjecture 1 says that these operators actually act non-trivially only in the subspace of cyclic invariant states. Conjecture 2 shows that these operators can be (somewhat unexpectedly) expressed in terms of the non-diagonal components of the total spin operator. Conjecture 3 states that they span in $B$- or $C$-type Lie algebra, according to the parity of the number of sites of the chain. Conjecture 4 asserts that the symmetry algebra of the model is actually wider and extended by a central element contained in the third component of the total spin operator.
In our previous study devoted to the Motzkin chain [P-25] the result analogous to Theorem 2 was formulated as a conjecture. In fact, the proof we give here for Theorem 2 can be directly generalized to the spin-$1$ case of Motzkin chain. This is intimately related to fact that the lowering and raising operators have essentially the same structure for both models.
Conjecture 1 can also be formulated in the context of Motzkin chain. It is not however clear how it can be proven in either case. Concerning Conjecture 2, it seems plausible that it can be proven by suitably applying the technique exposed in [ZF-11]. Conjectures 3 and 4 are direct analogues of the corresponding conjectures made in the context of Motzkin chain in [P-25]. We have observed that the central element for the Fredkin chain with even number of sites appears to be related to the third component of total spin operator by exactly the same relation as was found in the context of the Motzkin chain. In both cases the symmetry algebra is the $C$-type Lie algebra of the same rank. It would be interesting to understand group-theoretical backgrounds of these results. In addition, the coefficients $\gamma_{i}$ and $\alpha_{i}$ appearing in Conjecture 2 and Conjecture 4, respectively, certainly have some combinatorial interpretations that can also be a subject of further study.
@@ Acknowledgments
The author is indebted for stimulating discussions to Nikolai M. Bogoliubov, Vladimir E. Korepin, and Vitaly O. Tarasov. The author also thanks anonymous referees for useful remarks.